### v2

# Isomorphism in Union-Closed Sets

Abstract
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be pure union-closed families of sets. A map $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ is called an isomorphism if it is bijective and preserves unions: $h(A\cup B)=h(A)\cup h(B)$ for all $A,B\in\mathcal{K}_{1}$. We show that for any such isomorphism $h$, there exists a corresponding hyperisomorphism $H:\bigcup\mathcal{K}_{1}\to\bigcup\mathcal{K}_{2}$ such that $h(A)=H(A)$ for all $A\in\mathcal{K}_{1}$. This result provides a new framework for studying the structural properties of union-closed families and may offer insights into the union-closed set conjecture.
Keywords: hyperisomorphism, isomorphism, redundant element, union-closed sets, union-closed set conjecture
1 Introduction
For an integer $n\in\mathbb{N}$, we define the set $[n]$ as $\{k\in\mathbb{N}\mid k\leq n\}$. A family $\mathcal{K}\subseteq 2^{[n]}$ is called union-closed if for all $A,B\in\mathcal{K}$, it holds that $A\cup B\in\mathcal{K}$. In 1979, Péter Frankl [1] proposed a deceptively simple conjecture that has remained unresolved for over four decades, despite significant efforts and noteworthy progress. The conjecture is stated as follows:
Conjecture 1.1.
For every non-empty union-closed family of sets, there exists an element that belongs to at least half of the sets in the family.
In a previous paper [2], we proved a weaker version of this conjecture, which is summarized in the following theorem:
Theorem 1.1.
Let $\mathcal{K}\subseteq 2^{[n]}$ be a union-closed family of sets such that for all $A,B\in\mathcal{K}$, we have $A^{c}\cup B\in\mathcal{K}$, where $A^{c}=\bigcup\mathcal{K}\setminus A$. Then, there exists an element that belongs to at least half of the sets in the family.
In this article, we investigate the structural properties of isomorphisms between union-closed families of sets. Specifically, we prove in Theorem 3.1 that for every isomorphism $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ between two pure union-closed families of sets, there exists a corresponding hyperisomorphism $H:\bigcup\mathcal{K}_{1}\to\bigcup\mathcal{K}_{2}$ such that $h(A)=\{H(a)\mid a\in A\}$ for all $A\in\mathcal{K}_{1}$.
Definition 1.1 (Homomorphism).
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be union-closed families of sets. A mapping $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ is called a homomorphism if, for all $A_{1},A_{2}\in\mathcal{K}_{1}$, the following property holds:
$h(A_{1}\cup A_{2})=h(A_{1})\cup h(A_{2}).$
Definition 1.2 (Isomorphism).
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be union-closed families of sets. A homomorphism $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ is called an isomorphism if it is bijective.
Lemma 1.1.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two union-closed families, and $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ be a homomorphism between them. If $A_{1},A_{2}\in\mathcal{K}_{1}$ such that $A_{1}\subseteq A_{2}$, then:
$h(A_{1})\subseteq h(A_{2}).$
Proof.
By the homomorphism property, we have:
$h(A_{2})=h(A_{1}\cup A_{2})=h(A_{1})\cup h(A_{2}).$
Thus, $h(A_{1})\subseteq h(A_{2})$. ∎
Corollary 1.1.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two union-closed families, and $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ be an isomorphism between them. If $A_{1},A_{2}\in\mathcal{K}_{1}$ such that $A_{1}\subset A_{2}$, then:
$h(A_{1})\subset h(A_{2}).$
Lemma 1.2.
Let $\mathcal{K}$ be a union-closed family of set, and $h:\mathcal{K}\to 2^{[n]}$ be a homomorphism. Then the image of $\mathcal{K}$ under $h$ forms a union-closed family of sets.
Proof.
Let $B_{1},B_{2}\in h(\mathcal{K})$. Then there exist $A_{1},A_{2}\in\mathcal{K}$ such that $B_{1}=h(A_{1})$ and $B_{2}=h(A_{2})$. Therefore:
$B_{1}\cup B_{2}=h(A_{1})\cup h(A_{2})=h(A_{1}\cup A_{2}).$
Thus, $B_{1}\cup B_{2}\in h(\mathcal{K})$, proving that $h(\mathcal{K})$ is union-closed. ∎
Corollary 1.2.
Let $\mathcal{K}$ be a union-closed family of set, and $h:\mathcal{K}\to 2^{[n]}$ be an injective homomorphism. Then $h:\mathcal{K}\to h(\mathcal{K})$ is an isomorphism.
Definition 1.3 (Redundant Element).
Let $\mathcal{K}\subseteq 2^{[n]}$. An element $z\in\bigcup\mathcal{K}$ is called redundant, if removing $z$ from every set in $\mathcal{K}$ does not change the size of the collection. Specifically, we define:
$\mathcal{K}^{\setminus z}=\{X\setminus\{z\}\mid X\in\mathcal{K}\},$
where $|\mathcal{K}|=|\mathcal{K}^{\setminus z}|$. The collection $\mathcal{K}^{\setminus z}$ is called the reduced collection.
Definition 1.4 (Pure Collection).
A collection $\mathcal{K}\subseteq 2^{[n]}$ is called pure if it does not have any redundant element.
Let $\mathcal{K}$ be a collection of sets. The Purified Collection of $\mathcal{K}$, denoted by $\mathcal{K}^{*}$, is constructed by iteratively removing all redundant elements from $\mathcal{K}$. Although this concept is not yet rigorously defined, we demonstrate that all such purified collections are isomorphic. Furthermore, if $\mathcal{K}$ is already a pure collection of sets, then $\mathcal{K}=\mathcal{K}^{*}$.
Lemma 1.3.
Let $\mathcal{K}$ be a union-closed family of sets, and $z$ be a redundant element of $\mathcal{K}$. Then there exists an isomorphism between $\mathcal{K}$ and $\mathcal{K}^{\setminus z}$.
Proof.
Define $h:\mathcal{K}\to\mathcal{K}^{\setminus z}$ by $h(A)=A\setminus\{z\}$. Let $A,B\in\mathcal{K}$. Then:
$h(A\cup B)=(A\cup B)\setminus\{z\}=(A\setminus\{z\})\cup(B\setminus\{z\})=h(A)\cup h(B).$
This shows that $h$ preserves the union operation, so $h$ is a homomorphism. It is clear that $h$ is surjective because every element of $\mathcal{K}^{\setminus z}$ is of the form $A\setminus\{z\}$ for some $A\in\mathcal{K}$. To show that $h$ is injective, suppose $h(A)=h(B)$ for some $A,B\in\mathcal{K}$. This implies:
$A\setminus\{z\}=B\setminus\{z\}.$
If $A\neq B$, removing $z$ would change the size of $\mathcal{K}$, contradicting the assumption that $z$ is a redundant element. Thus, $A=B$, which implies that $h$ is injective. Since $h$ is both a homomorphism and a bijection, it follows that $h$ is an isomorphism between $\mathcal{K}$ and $\mathcal{K}^{\setminus z}$. ∎
Corollary 1.3.
For any union-closed family $\mathcal{K}$, there exists an isomorphism $h:\mathcal{K}\to\mathcal{K}^{*}$.
2 Cardinality Theorem
In this section, we present and prove a key theorem that forms the foundation for establishing our main result. Before stating the Cardinality Theorem, we introduce necessary terminology and prove supporting lemmas.
Definition 2.1.
Let $\mathcal{K}\subseteq 2^{[n]}$ be a poset under set inclusion. An element $X\in\mathcal{K}$ is called minimal if it has no proper subset in $\mathcal{K}$. The set of all minimal elements in $\mathcal{K}$ is denoted by $\mathcal{K}^{\bot}$ and is defined as:
$\mathcal{K}^{\bot}=\{X\in\mathcal{K}\mid\nexists A\in\mathcal{K}\text{ such that }A\subset X\}.$
Figure 1: Two lattice diagrams. The red nodes belong to $\mathcal{K}^{\bot}$.
Lemma 2.1.
Let $\mathcal{K}$ be a union-closed family, and $X\in\mathcal{K}^{\bot}$. Then $\mathcal{K}\setminus\{X\}$ is also union-closed.
Proof.
Let $X$ be a minimal element of $\mathcal{K}$, and consider any $A,B\in\mathcal{K}\setminus\{X\}$. Since $X$ is a minimal element, neither $A$ nor $B$ can be a proper subset of $X$. Thus, $A\cup B\neq X$. Furthermore, because $\mathcal{K}$ is union-closed, $A\cup B\in\mathcal{K}$. As $A\cup B\neq X$, we conclude that $A\cup B\in\mathcal{K}\setminus\{X\}$. Therefore, $\mathcal{K}\setminus\{X\}$ is union-closed, as required. ∎
Lemma 2.2.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be union-closed families, and $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ be an isomorphism between them. If $X$ is a minimal element of $\mathcal{K}_{1}$, then $h(X)$ is a minimal element of $\mathcal{K}_{2}$.
Proof.
Suppose, for the sake of contradiction, that $B\subset h(X)$ for some $B\in\mathcal{K}_{2}$. Then, because $h$ is surjective, there exists $A\in\mathcal{K}_{1}$ such that $h(A)=B$ and $A\neq X$. Now consider the set $h(A\cup X)$:
$h(A\cup X)=h(A)\cup h(X)=B\cup h(X)=h(X).$
Since $h$ is injective, it follows that $A\cup X=X$. This implies $A\subset X$, contradicting the minimality of $X$ in $\mathcal{K}_{1}$. Therefore, $h(X)$ is minimal in $\mathcal{K}_{2}$. ∎
Lemma 2.3.
Let $\mathcal{K}$ be a pure union-closed family of sets. If $|\mathcal{K}^{\bot}|=1$, then $\mathcal{K}^{\bot}=\{\varnothing\}$.
Proof.
By assumption, there exists exactly one set, denoted by $A$, that belongs to $\mathcal{K}^{\bot}$. It is straightforward to verify that every maximal chain in $\mathcal{K}$ must start from a member of $\mathcal{K}^{\bot}$. In this case, every maximal chain can be represented as:
$A\subset C_{1}\subset C_{2}\subset\cdots\subset\bigcup\mathcal{K}.$
Since every element $X\in\mathcal{K}$ lies on a maximal chain, it follows that $A\subset X$. If $A$ contained any elements, those elements would be redundant, as $A$ is a subset of all elements in $\mathcal{K}$. However, because $\mathcal{K}$ is a pure collection of sets, $A$ must be the empty set. ∎
Theorem 2.1 (Cardinality).
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two pure union-closed families of sets. If there exists an isomorphism $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$, then for every $A\in\mathcal{K}_{1}$, we have:
$|A|=|h(A)|.$
Proof.
We prove the statement by induction.
Base Case: Consider $|\mathcal{K}_{1}|=|\mathcal{K}_{2}|=1$. In this case, $\mathcal{K}_{1}=\{A\}$ and $\mathcal{K}_{2}=\{B\}$ for some $A,B\in 2^{[n]}$. Both $A$ and $B$ must necessarily be empty sets. If $A$ or $B$ were not empty, their elements would be redundant, and removing those elements would not affect the cardinality of $\mathcal{K}_{1}$ or $\mathcal{K}_{2}$. Thus, for the base case, we have $|\bigcup\mathcal{K}_{1}|=|\bigcup\mathcal{K}_{2}|=0$, as desired.
Inductive Step: Assume that the theorem holds for all pure union-closed families with cardinality strictly less than $n$. Now, consider $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ such that $|\mathcal{K}_{1}|=|\mathcal{K}_{2}|=n$, and let $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ be an isomorphism. Let $X\in\mathcal{K}_{1}^{\bot}$. By Lemma 2.2, we have $h(X)=Y$ where $Y\in\mathcal{K}_{2}^{\bot}$. Consider the restriction $h:\mathcal{K}_{1}\setminus\{X\}\to\mathcal{K}_{2}\setminus\{Y\}$. By Lemma 2.1, this restricted mapping remains an isomorphism.
Case 1:
Suppose $\mathcal{K}_{1}\setminus\{X\}$ and $\mathcal{K}_{2}\setminus\{Y\}$ are pure collections. By the induction hypothesis, for all $A\in\mathcal{K}_{1}\setminus\{X\}$, we have $|A|=|h(A)|$. If $|\mathcal{K}_{1}^{\bot}|=1$, then by Lemma 2.3, $X=\varnothing$ and $Y=\varnothing$ as well. Consequently, for all $A\in\mathcal{K}_{1}$, we conclude $|A|=|h(A)|$.
On the other hand, if $|\mathcal{K}_{1}^{\bot}|>1$, let $Z\in\mathcal{K}_{1}^{\bot}$ such that $Z\neq X$. If $\mathcal{K}_{1}\setminus\{Z\}$ and $\mathcal{K}_{2}\setminus\{h(Z)\}$ are pure collections, then by the induction hypothesis, we again obtain $|X|=|Y|$, completing the proof for this case.
Thus, the theorem is proved for the scenario where removing one element of $\mathcal{K}_{1}^{\bot}$ and its corresponding map results in pure collections. Next, we consider the case where there exists at least one element in $\mathcal{K}_{1}^{\bot}$ or $\mathcal{K}_{2}^{\bot}$ such that its removal results in a non-pure collection.
Case 2:
Without loss of generality, suppose $\mathcal{K}_{1}\setminus\{X\}$ is not pure and contains a redundant element, denoted by $z$. To proceed, we establish three key facts:
  1. 1.
    $z\in C$ for all $C\in\mathcal{K}_{1}\setminus\{X\}$.
  2. 2.
    $\mathcal{K}_{1}\setminus\{X\}$ has at most one redundant element.
  3. 3.
    If $\mathcal{K}_{1}\setminus\{X\}$ has a redundant element, then $\mathcal{K}_{2}\setminus\{Y\}$ also has a redundant element.
First, consider $z$ as a redundant element of $\mathcal{K}_{1}\setminus\{X\}$ but not of $\mathcal{K}_{1}$. This implies there exists exactly one element $A_{X}\in\mathcal{K}_{1}\setminus\{X\}$ such that $A_{X}\setminus\{z\}=X$, and no distinct $A_{1},A_{2}\in\mathcal{K}_{1}\setminus\{X\}$ satisfy $A_{1}\setminus\{z\}=A_{2}$. Now, for the sake of contradiction, suppose there exists $B\in\mathcal{K}_{1}\setminus\{X\}$ such that $z\notin B$. Then $z\notin B\cup X$. Let $A_{1}=A_{X}\cup B$ and $A_{2}=B\cup X$. It is straightforward to verify that $A_{1}\setminus\{z\}=A_{2}$, leading to a contradiction. Thus, if $z$ is a redundant element of $\mathcal{K}_{1}\setminus\{X\}$, it must belong to all its members.
Next, suppose $t$ is another redundant element of $\mathcal{K}_{1}\setminus\{X\}$. From the previous fact, we know that $t$ must belong to all members of $\mathcal{K}_{1}\setminus\{X\}$ but not to $X$. However, $t$ cannot satisfy this condition because $A_{X}\setminus\{z\}=X$. Thus no additional redundant element can exist. Therefore, $\mathcal{K}_{1}\setminus\{X\}$ can contain at most one redundant element.
For the third fact, let $z$ be the redundant element of $\mathcal{K}_{1}\setminus\{X\}$ and $A_{X}=X\cup\{z\}$. Define $R=h(A_{X})\setminus Y$, which is non-empty by Corollary 1.1. We show that $R$ is a subset of all members of $\mathcal{K}_{2}\setminus\{Y\}$. For the sake of contradiction, suppose there exists $D\in\mathcal{K}_{2}\setminus\{Y\}$ such that $R\not\subset D$. By fact 1, we know $z\in h^{-1}(D)$. Therefore:
$h^{-1}(D)\cup X=h^{-1}(D)\cup A_{X}\implies D\cup Y=D\cup h(A_{X}).$
This leads to a contradiction, since $R$ is not a subset of $D\cup Y$, but it is a subset of $D\cup h(A_{X})$. Thus, $R$ must be a subset of all members of $\mathcal{K}_{2}\setminus\{Y\}$. Therefore, the members of $R$ are redundant. As previously proven, $\mathcal{K}_{2}\setminus\{Y\}$ has at most one redundant element. Hence, $|R|=1$, and we denote $R=\{r\}$.
The mapping $h^{*}:{(\mathcal{K}_{1}\setminus\{X\})}^{*}\to{(\mathcal{K}_{2}\setminus\{Y\})}^{*}$, defined as $h^{*}(A\setminus\{z\})=h(A)\setminus\{r\}$ where $A\in\mathcal{K}_{1}\setminus\{X\}$, is an isomorphism between two pure union-closed families of sets. Furthermore, the cardinalities satisfy $|{(\mathcal{K}_{1}\setminus\{X\})}^{*}|=|{(\mathcal{K}_{2}\setminus\{Y\})}^{*}|=n-1$. By the induction hypothesis, for all $A\setminus\{z\}\in{(\mathcal{K}_{1}\setminus\{X\})}^{*}$, we have:
$|A\setminus\{z\}|=|h^{*}(A\setminus\{z\})|=|h(A)\setminus\{r\}|.$
Since for all $A\in\mathcal{K}_{1}\setminus\{X\}$, we have $z\in A$ and $r\in h(A)$, it follows that $|A|=|h(A)|$.
From fact 1 and fact 3, there exists a set $A_{X}\in\mathcal{K}_{1}$ such that $z\in A_{X}$ but $z\not\in X$, and $r\in h(A_{X})$ but $r\not\in Y$. This implies:
$|X|=|A_{X}|-1=|h(A_{X})|-1=|Y|.$
Thus, for all $A\in\mathcal{K}_{1}$, it follows that $|A|=|h(A)|$.
We conclude this section with two important corollaries that aid in proving the main theorem of this paper.
Corollary 2.1.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two pure union-closed families of sets. If there exists an isomorphism $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$, then:
$\left|\bigcup\mathcal{K}_{1}\right|=\left|\bigcup\mathcal{K}_{2}\right|.$
Corollary 2.2.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two pure union-closed families of sets. If there exists an isomorphism $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$, then for any $A,B\in\mathcal{K}_{1}$, the following properties hold:
  1. 1.
    $|A\cup B|=|h(A)\cup h(B)|$.
  2. 2.
    $|A\cap B|=|h(A)\cap h(B)|$.
  3. 3.
    $|A\setminus B|=|h(A)\setminus h(B)|$.
  4. 4.
    $|A^{c}|=|h{(A)}^{c}|$,
where $A^{c}=\bigcup\mathcal{K}_{1}\setminus A$ and $h{(A)}^{c}={\bigcup\mathcal{K}_{2}}\setminus h(A)$.
3 Hyperisomorphism
In this section, we prove that the structure of a pure union-closed family of sets is entirely preserved under an isomorphism.
Definition 3.1 (Hyperisomorphism).
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two union-closed families of sets. A bijective mapping $H:\bigcup\mathcal{K}_{1}\to\bigcup\mathcal{K}_{2}$ is called a hyperisomorphism if the induced mapping $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$, defined by $h(A)=\{H(a)\mid a\in A\}$, is an isomorphism.
To establish our main theorem, we require the following lemmas. First, we introduce the notation $\mathcal{K}^{i}$, defined as:
$\mathcal{K}^{i}=\{A\in\mathcal{K}\mid i\in A\}.$
Lemma 3.1.
Let $\mathcal{K}$ be a pure union-closed family of sets. Then, for all $i,j\in\bigcup\mathcal{K}$, the following equivalence holds:
$i=j\iff\mathcal{K}^{i}=\mathcal{K}^{j}.$
Proof.
Suppose $i=j$. By definition, it follows that $\mathcal{K}^{i}=\mathcal{K}^{j}$. Conversely, assume $\mathcal{K}^{i}=\mathcal{K}^{j}$, but $i\neq j$. This would imply that $i$ and $j$ are redundant elements in $\mathcal{K}$, contradicting the purity of $\mathcal{K}$. Hence, it must be the case that $i=j$. ∎
Lemma 3.2.
Let $\mathcal{K}$ be a union-closed family of sets. Then, for any distinct elements $i,j\in\bigcup\mathcal{K}$, the following inequality holds:
$\left|\left(\bigcap\mathcal{K}^{i}\right)\cup\left(\bigcap\mathcal{K}^{j}\right)\right|\geq 2.$
Proof.
This follows since $i$ and $j$, being distinct, are both in the union. ∎
Corollary 3.1.
Let $\mathcal{K}$ be a union-closed family of sets. Then, for any distinct elements $\{a_{i}\}_{i=1}^{n}\subseteq\bigcup\mathcal{K}$, the following inequality holds:
$\left|\bigcup_{i=1}^{n}\bigcap\mathcal{K}^{a_{i}}\right|\geq n.$
Lemma 3.3.
Let $\mathcal{K}$ be a union-closed family of sets, and $i$ be an arbitrary element of $\bigcup\mathcal{K}$. If $\bigcap\mathcal{K}_{1}^{i}=\{a_{1},\dots,a_{n}\}$, then:
$\left|\bigcup_{j=1}^{n}\bigcap\mathcal{K}_{1}^{a_{j}}\right|=n.$
Proof.
Let $a$ be an arbitrary element of $\bigcap\mathcal{K}^{i}$. Since $\mathcal{K}^{i}\subseteq\mathcal{K}^{a}$, it follows that $\bigcap\mathcal{K}^{a}\subseteq\bigcap\mathcal{K}^{i}$. Consequently, we have:
$\bigcup_{j=1}^{n}\bigcap\mathcal{K}^{a_{j}}\subseteq\bigcap\mathcal{K}^{i},$
which implies:
$\left|\bigcup_{j=1}^{n}\bigcap\mathcal{K}_{1}^{a_{j}}\right|\leq n.$
Applying Corollary 3.1, we conclude that:
$\left|\bigcup_{j=1}^{n}\bigcap\mathcal{K}_{1}^{a_{j}}\right|=n.$
This completes the proof. ∎
Lemma 3.4.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two pure union-closed families of sets, and $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$ be an isomorphism between them. Then, for each $i\in\bigcup\mathcal{K}_{1}$, there exists a unique $j\in\bigcup\mathcal{K}_{2}$ such that:
$\mathcal{K}_{2}^{j}=h(\mathcal{K}_{1}^{i}),$
where $h(\mathcal{K}_{1}^{i})=\{h(A)\mid A\in\mathcal{K}_{1}^{i}\}$.
Proof.
Let $i\in\bigcup\mathcal{K}_{1}$. Suppose $\bigcap\mathcal{K}_{1}^{i}=\{a_{1},\dots,a_{n}\}$. By Lemma 3.3, we have:
$\left|\bigcup_{j=1}^{n}\bigcap\mathcal{K}_{1}^{a_{j}}\right|=n.$
By Corollary 2.2, we can assume $\bigcap h(\mathcal{K}_{1}^{i})=\{b_{1},\dots,b_{n}\}$.
Since $h(\mathcal{K}_{1}^{i})\subseteq\mathcal{K}_{2}^{b}$ for each $b\in\bigcap h(\mathcal{K}_{1}^{i})$, it suffices to show that for exactly one $b_{i}\in\bigcap h(\mathcal{K}_{1}^{i})$ we have:
$|h(\mathcal{K}_{1}^{i})|=|\mathcal{K}_{2}^{b_{i}}|.$
Assume, for the sake of contradiction, that for all $b\in\bigcap h(\mathcal{K}_{1}^{i})$, we have $|h(\mathcal{K}_{1}^{i})|<|\mathcal{K}_{2}^{b}|$. This implies $i\notin\bigcap h^{-1}(\mathcal{K}_{2}^{b})$, and thus:
$\bigcap h^{-1}(\mathcal{K}_{2}^{b})\subset\bigcap\mathcal{K}_{1}^{i}.$
Consequently:
$\bigcup_{j=1}^{n}\bigcap h^{-1}(\mathcal{K}_{2}^{b_{j}})\subset\bigcap\mathcal{K}_{1}^{i},$
which implies:
$\left|\bigcup_{j=1}^{n}\bigcap h^{-1}(\mathcal{K}_{2}^{b_{j}})\right|<n.$
Then by Corollary 2.2, we then have:
$\left|\bigcup_{j=1}^{n}\bigcap\mathcal{K}_{2}^{b_{j}}\right|<n.$
However, Corollary 3.1 ensures that:
$\left|\bigcup_{j=1}^{n}\bigcap\mathcal{K}_{2}^{b_{j}}\right|\geq n,$
which is a contradiction. Therefore, by Lemma 3.1, there exists exactly one $j\in\bigcap h(\mathcal{K}_{1}^{i})$ such that:
$\mathcal{K}_{2}^{j}=h(\mathcal{K}_{1}^{i}).$
This completes the proof. ∎
Theorem 3.1.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two pure union-closed families of sets. For every isomorphism $h:\mathcal{K}_{1}\to\mathcal{K}_{2}$, there exists a hyperisomorphism $H:\bigcup\mathcal{K}_{1}\to\bigcup\mathcal{K}_{2}$ such that:
$h(A)=\{H(a)\mid a\in A\}\quad\text{for all }A\in\mathcal{K}_{1}.$
Proof.
Let $i\in\bigcup\mathcal{K}_{1}$. By Lemma 3.4, there exists a corresponding $j\in\bigcup\mathcal{K}_{2}$ which allows us to define $H$. To show that $H$ is bijective, suppose $H(a)=H(b)$ for some $a,b\in\bigcup\mathcal{K}_{1}$. This implies that:
$\mathcal{K}_{1}^{a}=\mathcal{K}_{1}^{b}.$
Using Lemma 3.1, we conclude that $a=b$, establishing that $H$ is injective. Furthermore, by Corollary 2.1, $H$ is clearly surjective. Thus, $H$ is bijective. Next, let $A\in\mathcal{K}_{1}$ and $i\in A$. To complete the proof, we need to show that $H(i)\in h(A)$. By Lemma 3.4, we have
$h(A)\in\mathcal{K}_{2}^{H(i)},$
which satisfies the required condition. ∎